### v1

Abstract
One of the most dominant energy budget of the Universe is Dark Energy, which remains enigmatic since the claim of its existence from the observation of late-time cosmic acceleration. We propose a new way of inferring this by measuring the aging of the Universe using only gravitational wave (GW) signals from coalescing binary compact objects of any masses. We show that the aging of the Universe will lead to a change in the observed chirp mass of the GW sources inferred from different stages of its coalesces by monitoring a coherent source from two far-apart GW frequencies for a few years. We show that coordinated GW detectors which can reach a relative uncertainty on chirp mass measurement of about $10^{-9}$, can measure a tiny departure of about $2\%$ from the dark energy equation of state parameter $w_{0}=-1$ and its variation with cosmic time by using stellar origin binary black holes up to redshift $z=5$ in 10 years of observation time without using any external calibrator. If the next generation of GW detectors can achieve this precision, then it can open a new window to discover the fundamental nature of dark energy.
keywords:
Cosmology, Dark Energy, Gravitational Waves.
1 Introduction
The evidence of an accelerating Universe detected from cosmological observations has indicated the presence of about $75\%$ of the energy budget in dark energy [1, 2, 3, 4, 5, 6]. The nature of dark energy remains to be unknown for several decades and multiple experiments are ongoing to understand it. Recently, the observation of Baryon Acoustic Oscillation (BAO) from DESI in combination with supernovae and CMB has hinted towards a redshift-dependent model of dark energy density, which can have a profound implication if the results are not influenced by systematic [7]. We show in this paper that by using only Gravitational waves (GWs) signals from coalescing binary compact objects, we can reach an unprecedented precision to measure the nature of dark energy without requiring any calibration or astrophysical assumption on the source.
GWs– a new cosmic messenger– can probe the Universe up to high redshift from the current generation (such as Advanced-LIGO (Hanford, Livingston, and Aundha) [8, 9], Virgo [10], KAGRA [11]) and next generation GW detectors (such as Cosmic Explorer [12], Einstein Telescope[13, 14], LISA [15], and concepts such as LGWA [16]). As the signal emitted from coalescing binary compact objects is extremely well modeled in the General Theory of Relativity (GR), it makes it possible to accurately calculate the change in the GW frequency and amplitude with time depending on the source properties. As a result, the GW sources are both standard clocks and standard sirens that make it possible to measure the chirp mass and luminosity distance to the GW sources accurately [17, 18].
In this work, we show that by using two salient aspects of GW signal, namely (i) we can measure both redshifted chirp-mass and luminosity distance to the GW sources with reasonable accuracy and (ii) the period of observation of the GW signals from a single source can be measured in phase for a year(s) with the aid of multi-band GW observatories operating at different GW frequencies, we can measure the nature of dark energy without requiring any additional information. The long-period observation of GW sources makes it possible to measure the change in the expansion rate of the Universe from its aging as a function of luminosity distance, and by exploiting these, we can map the expansion history up to the high redshift Universe.
2 Cosmology from the aging of the Universe
[…]
$H(z)=H_{0}\bigg{[}\Omega_{\rm dm}(z)+\Omega_{\rm b}(z)+\Omega_{\rm de}(z)+\Omega_{\rm rad}(z)\bigg{]}^{1/2},$ (1)
where $H_{0}$ is the Hubble constant which captures the expansion rate today in the Universe, and $\Omega_{i}(z)\equiv\rho(z)/\rho_{c}$ denotes the energy density in the $i^{th}$ constituent of the Universe with respect to the critical energy density $\rho^{2}_{c}\equiv 3H_{0}^{2}/8\pi G$. As per our current understanding, about $95\%$ of today’s Universe is made up of dark energy (denoted by $\Omega_{\rm de}$) and cold dark matter (denoted by $\Omega_{\rm dm}$), which are still unknown to us, and the remaining $5\%$ contribution comes from baryons and radiation (denoted by $\Omega_{\rm b}$, $\Omega_{\rm rad}$, ) [19].
Figure 1: The relative variation of redshift due to different models of cosmic expansion history is shown between two epochs separated by $\Delta t=5$ years. The fiducial cosmological model is considered for the LCDM model with Planck-2018 cosmological parameters and $H_{0}=70$ km/s/Mpc.
However, these fractions change in the earlier epoch of the Universe, and measuring these fractions at different cosmic epochs enables us to understand the nature of these quantities primarily dark matter and dark energy, which are not well understood. The time-variation of dark energy is usually modeled by a vanilla parametric form of the dark energy EoS $w(a)=w_{0}+w_{a}(1-a(t))$ where $w_{0}$ is a constant and $w_{a}$ captures the time evolution of the dark energy EoS in terms of the cosmological scale factor $a(t)$ [20, 21]. The best-fit cosmological model supports $w_{0}=-1;w_{a}=0$ (cosmological constant) [19]. However, the recent DESI observations have indicated the signature of redshift dependent cosmological constant showing a best-fit value of $w_{0}>-1;w_{a}<0$ [7]. Furthermore, the limitation of this vanilla parametric form should not be neglected [22].
Direct evidence of cosmic expansion can also be inferred from the change in the aging of the Universe. This effect reflects to a change in the cosmic redshift $z$ with time. Given that the relation between redshift and scale factor $z=a(t_{0})/a(t)-1$, we can write the change in the redshift of a source with respect to the observed time $t$ as
$\frac{d\ln(1+z)}{dt}=H_{0}-\frac{H(z)}{(1+z)}.$ (2)
This implies that the fractional change in the redshift to the source depends on the difference between the Hubble parameter today (at $z=0$) and at a redshift $z$. An observation that can measure the change in the redshift of a source between an observation time separated by $\Delta t$ precisely, can probe the expansion history. In Fig. 1, we show the variation in redshift for different cosmological models for $\Delta t=5$ years.
3 Gravitational wave as a probe to cosmic aging
GW signal in terms of frequency $f$ from a coalescing binary of masses $m_{1}$ and $m_{2}$ situated at a luminosity distance $d_{L}$ can be written in the inspiral phase in terms of the source-frame chirp mass $\mathcal{M}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ as [23, 24]
$h_{+,\times}(f,\hat{n})=\sqrt{\frac{5}{96}}\frac{G^{5/6}\mathcal{M}^{2}(f\mathcal{M})^{-7/6}}{c^{3/2}\pi^{2/3}d_{L}}{\Theta}_{+,\times}(\hat{L}.\hat{n}),$ (3)
where ${\Theta}_{+,\times}(\hat{L}.\hat{n})$ is the factor that captures the projection between the orbital angular momentum $\hat{L}$ and the direction to the line of sight to the observer $\hat{n}$ for the plus ($h_{+}$) and cross-polarization ($h_{\times}$) states of the GW signal. A key property of coalescing binary objects is the change in the frequency of the signal with time [23, 24]
$\frac{df}{dt}=\frac{96\pi^{8/3}}{5c^{5}}(G\mathcal{M})^{5/3}f^{11/3}.$ (4)
This shows that the change in the frequency of the GW signal depends strongly on the chirp mass of the coalescing binaries and the instantaneous frequency of the GW signal. In an expanding Universe, the time gets dilated as $dt_{o}=dt(1+z)$ and frequency gets redshifted as $f_{o}=f/(1+z)$, which leads to the redshifted observed chirp mass $\mathcal{M}^{\rm det}=(1+z)\mathcal{M}$. This implies
$\frac{df_{o}}{dt_{o}}(\tau)=\frac{96\pi^{8/3}}{5c^{5}}(G(1+z(\tau))\mathcal{M})^{5/3}f_{o}^{11/3},$ (5)
Figure 2: This schematic diagram explains the key principle behind this technique to measure the expansion history of the Universe. A GW source emitting at a low frequency witnesses a different cosmic expansion history than when it is emitting at a later time at a high frequency due to the aging of the Universe. As a result, the inferred detector-frame chirp mass from low frequency $\mathcal{M}^{\rm det}_{\rm low}(\tau)=(1+z(\tau))\mathcal{M}$ will be different from the detector-frame chirp mass from high frequency $\mathcal{M}^{\rm det}_{\rm high}(\tau+\Delta t)=(1+z(\tau+\Delta t))\mathcal{M}$ due to the aging of the Universe.
where $\tau$ denotes a cosmic epoch. The above equation shows that the rate of evolution of the GW frequency in the early inspiral phase and the late inspiral (or in the merger/ringdown phase) will be happening at a different cosmic redshift. As a result, the corresponding chirping behavior (Eq. (5)) of the signal will be tracing a little different value of $z(\tau)$ due to the cosmic expansion (see the schematic diagram in Fig. 2).
So if we observe a coherent GW source detected at a low frequency during an early inspiral phase and also at a high frequency during the late inspiral phase, they will have a little different detector-frame chirp mass denoted by $\mathcal{M}^{\rm det}_{\rm low/high}(\tau)=(1+z(\tau))\mathcal{M}$. Their relative difference can be connected to cosmic expansion history as
$\Delta\mathcal{M}^{det}_{high,low}\equiv\frac{\mathcal{M}^{\rm det}_{\rm high}(\tau+\Delta t)}{\mathcal{M}^{\rm det}_{\rm low}(\tau)}-1$ $=\frac{z(\tau+\Delta t)-z(\tau)}{1+z(\tau)},$ (6)
$\Delta\mathcal{M}^{det}_{high,low}$ $=\Delta t\bigg{(}H_{0}-\frac{H(z(\tau))}{1+z(\tau)}\bigg{)},$
where, in the last equation, we have used Eq. (2) to show its relation with the change in the cosmic expansion rate. The value of $\Delta t$ for GW sources such as stellar origin BBHs can be a few years if a source is observed in their inspiral stage at frequencies $10^{-3}$ Hz and also at a later stage in their inspiral when they are emitting at $\sim 10^{2}$ Hz. Combining the relative chirp mass difference measurement with another well-measured quantity, the luminosity distance to the source that is inferred from the average of the signal from the complete observation period, we can infer the expansion history of the Universe and cosmological parameters without requiring any EM counterpart, small sky-localization error, additional galaxy-catalog, or making any assumption on the GW mass population. A few key advantages of this method are:
  1. 1.
    Due to the coherent nature of the GW signal and its robust predictability from GR, the calibration of the measured luminosity distance and chirp mass is not required.
  2. 2.
    This method allows to study of cosmology with only GW sources without requiring an independent redshift measurement or better sky localization of the sources.
  3. 3.
    The method can robustly probe expansion rate up to a very high redshift without any assumption on the astrophysical population of sources.
Figure 3: We show the uncertainty in the inference of the expansion history for the fiducial LCDM model of cosmology up to $z=5$ using BBHs integrated over a period of 10 years for a chirp mass error of $10^{-9}$ and relative luminosity distance error of $10^{-1}$ by observing sources in GW frequencies $\sim 10^{-2}$ Hz and $\sim 10^{2}$ Hz, which are separated by a time difference $\Delta t=5$ years.
4 Results
The measurement of the high redshift expansion rate of the Universe demonstrated in the last section is possible from coherent GW sources such as binary neutron stars (BNSs), neutron star-black holes (NSBHs), binary black holes (BBHs) of all masses, which can include black holes of both astrophysical and primordial origin. In this analysis, we only consider stellar origin BBHs as it is a population that can be detected up to high redshift and also we have an estimate of its merger rate from LIGO-Virgo-KAGRA collaboration observations [25]. We obtain the number of GW mergers at any redshift by integrating the merger rate density $R(z)$ of the GW sources of some masses $m_{1},m_{2}$ over redshift by multiplying over the comoving volume $dV/dz$ and operational time $T_{\rm obs}$ of GW detectors as
[…]
Figure 4: The figure shows the Cramer-Rao bound on the dark energy EoS $w_{0}=-0.8$ and $w_{a}=-2$ (left) and $w_{0}=-1.0$ and $w_{a}=0$ (right) for two different relative chirp mass errors (shown in red ($10^{-8}$) and blue ($10^{-9}$)), after combining 10 years of observation of stellar origin BBHs up to $z=5$.
We model the merger rate of stellar origin BBHs $R(z)$ assuming that the BBHs follow the Madau-Dickinson star formation rate [26]. The local merger rate $R(z=0)$ is taken as 20 $\rm{Gpc}^{-3}\,\rm{yr}^{-1}$ as per the third observation run of LVK Collaboration (GWTC-3) [25]. This gives us in total about $10^{5}$ BBHs per year across all redshifts. This number can increase by about a factor of 2-3 if, we include the upper bound of BBH merger rate from GWTC-3 [25], and also combine BNS and NSBH sources.
The measurability of the expansion rate using this technique from a single GW event depends on the fractional standard deviation in measuring the masses in the low and high frequencies as
$\Sigma_{\Delta_{M}}=\sqrt{\bigg{(}\frac{\sigma_{\mathcal{M_{\rm low}}}}{\mathcal{M_{\rm low}}}\bigg{)}^{2}+\bigg{(}\frac{\sigma_{\mathcal{M_{\rm high}}}}{\mathcal{M_{\rm high}}}\bigg{)}^{2}},$ (8)
where $\frac{\sigma_{\mathcal{M_{\rm low,high}}}}{\mathcal{M_{\rm low,high}}}$ are the relative error on the mass measurements from the low frequency and high frequency of GW data. To explore the measurability of this signal, we consider two scenarios for the relative error on the chirp mass namely, $\frac{\sigma_{\mathcal{M_{\rm low,high}}}}{\mathcal{M_{\rm low,high}}}=10^{-8}$ and $10^{-9}$ for two frequency range of the GW inspiral stage which is separated by $\Delta t=5$ years. For comparison, the expected relative uncertainty on the chirp mass from current proposals [13, 15, 12, 16], is of the order $10^{-5}-10^{-6}$. So, the uncertainty in the chirp mass required for this analysis is about $10^{2}-10^{3}$ better than the currently proposed experiments. Though this improvement in uncertainty is large, the feasibility of this with the advancement of GW detector technology can make groundbreaking discoveries as will be discussed later in the paper. In particular, from the measurement perspective, precise inference of $\dot{f}/f$ over a broad range of frequencies is required for this measurement. Along with precise chirp mass inference, the precision on luminosity distance should be high as well. The relative luminosity distance uncertainty from individual GW sources ($\Sigma_{d_{L}}\equiv[(\sigma^{\rm det}_{d_{L}})^{2}+(\sigma^{\rm wl}_{d_{L}})^{2}]^{0.5}/d_{L}$) will arise primarily from two sources, uncertainty due to detector noise $\sigma^{\rm det}_{d_{L}}$ and due to weak lensing $\sigma^{\rm wl}_{d_{L}}$. For a relative luminosity distance uncertainty of about $\sigma^{\rm det}_{d_{L}}/d_{L}=1\%$, the dominant source of uncertainty is due to weak lensing for GW sources at high redshift ($z>1$) [27]. The requirement on the luminosity distance uncertainty for this technique is not stringent in comparison to the expected distance uncertainly from the next-generation GW detectors [13, 15, 12, 16].
We estimate the error in the inference of time-dependent chirp mass $\Delta\mathcal{M^{\rm det}_{\rm high,low}}$ as a function of luminosity distance in Fig. 3 for the fiducial case of LCDM model ($w_{0}=-1$ and $w_{a}=0$) and $\frac{\sigma_{\mathcal{M_{\rm low,high}}}}{\mathcal{M_{\rm low,high}}}=10^{-9}$. The plot shows that by using only observations of chirp mass with luminosity distance, we can measure the expansion history for a typical value of the astrophysical population of BBHs using the technique proposed in this work. We further explore the prospect of this technique to measure the phenomenological parameters $w_{0},w_{a}$ which captures the dark energy EoS and its redshift evolution. In Fig. 4 we show the Cramér-Rao bound [28, 29] on the minimum error in the inference of the dark energy EoS parameters ( $w_{0},w_{a}$) using Fisher analysis. The red and blue contours show the expected uncertainty for two cases of chirp mass errors $10^{-8}$ and $10^{-9}$ respectively, with 10 years of observation period for two choices of the dark energy EoS parameters, (a) $w_{0}=-0.8$; $w_{a}=-2$, and (b) $w_{0}=-1$; $w_{a}=0$. The plot indicates that with about $10^{-8}$ relative error on chirp mass, we can reach about a $5$-$\sigma$ detection of dark energy EoS within 10 years of the observation period. However, this measurement can reach a percent level ($\sim 2\%$) measurement of dark energy EoS and its redshift evolution for a relative chirp mass uncertainty $10^{-9}$. As a result, it will make a definite discovery on the nature of dark energy using observational probes that are free from any external calibration. If the number of GW sources detectable per year in multiple GW bands is more/less than $10^{5}$, then the required uncertainty on chirp mass measurement for the same precision to measure dark energy EoS will scale as $10^{-9}\sqrt{N_{GW}/10^{5}}$. This method can also explore the Hubble constant and other cosmological parameters, but we restrict this analysis only to dark energy, as it is one of the unsolved problems in cosmology.
5 Discussion and Conclusion
In this work, we show that in the framework of GR, observations of the shift in the detector frame chirp mass of the GW sources in two different frequency bands and inference of the luminosity distance provide a new way to measure the expansion rate of the Universe without requiring the use of any EM observations nor leveraging on any astrophysical assumption of the GW source. Due to the advantage of the GW signal not being susceptible to contamination from foreground contamination such as dust, the measure can be possible up to high redshift, to the epoch of the first ever binary black hole in the Universe.
This new frontier of observational cosmology if realized from the GW detectors, can provide the best possible way to measure the high redshift Universe. However, such measurements will require control of several sources of errors both on the instrument side as well as on the waveform modeling side. This kind of measurement will require waveform accuracy in the inspiral and merger phase up to $10^{-11}$ to successfully infer cosmology from GW signals. The degeneracy between inclination angle and luminosity distance [18], and impact from weak lensing [27] are dominant sources of uncertainties on the luminosity distance side. On the chirp mass inference side, the chirp mass inference can be impacted by environmental effects. As we are only focusing on stellar origin BBHs, these black holes are unlikely to have accretion discs or not be embedded only near a galactic center. As a result, the dephasing of the GW emission due to any local astrophysical effects is negligible.
As we are at an exciting time of designing the next generation of GW detectors, appropriate planning for observational strategies and advanced detector sensitivities can make a paradigm shift in observational cosmology using this technique. This technique brings a way to measure the nature of dark energy which is not limited by astrophysical modeling uncertainties and can be achieved by precise measurement from GW detectors operational from mili-Hertz to hecto-Hertz range. Beyond dark energy, this technique makes it also possible to independently map the expansion rate at high redshift, which can discover new components in the Universe that may be dominant only at high redshift. In summary, GW detector development which can make robust and precise measurements of the chirp mass, can open a new window to fundamental physics. This technique makes GW a self-sufficient probe to explore cosmology without depending on any EM observations.
[…]
Acknowledgements This work is part of the ⟨data|theory⟩ Universe-Lab, funded by TIFR and the Department of Atomic Energy, Government of India. The author acknowledges the use of computer cluster of the ⟨data|theory⟩ Universe-Lab and using Astropy [30, 31], Giant-Triangle-Confusogram [32], IPython [33], Matplotlib [34], NumPy [35], and SciPy [36] in this analysis.